OpenStudy (solomonzelman):

For real numbers a, b, and c, ab = 1.5, bc = 6, and ac = 25. So abc = (saw that online)

OpenStudy (solomonzelman):

I did the following I am not sure though

hartnn (hartnn):

multiply all the 3, you will have a^2b^2c^2 then just take square root

OpenStudy (solomonzelman):

(ab)(bc)(ac) = (1.5)(6)(25) aabbcc = 225 a^2b^2c^2 = 225 (abc)^2 = 225 abc = 15. is this right?

hartnn (hartnn):

yup

hartnn (hartnn):

correct and fastest method

OpenStudy (solomonzelman):

Good! Thank you!

hartnn (hartnn):

welcome ^_^

OpenStudy (solomonzelman):

I was confused b/c it seemed hard...